OpenStudy (anonymous):

In a triangle the ratio of the length of the sides are 19:11:8. Find the length of the three side if the area is 576cm^2.

OpenStudy (anonymous):

check ur question the lengths of the triangle can not be in the ratio of 19:11:8 as 8+11=19

OpenStudy (anonymous):

I just made that up. The real given is 17:10:9

OpenStudy (anonymous):

lol why would you change the question? XD

OpenStudy (anonymous):

Soo, i'll do it myself once i know the formula and pellet. xD

OpenStudy (anonymous):

assume the lengths to be 17x,10x,9x we have the semiperimetr as 36x/2= 18x and then proceed using Heron's formula for the area of the triangle

OpenStudy (anonymous):

Soo, x = 4?

OpenStudy (anonymous):

given sqrt(18x *(18x-17x)*(18x-10x)*(18x-9x)) =576 sqrt (18*8*9) x^2 =576 or 36x^2=576 or x^2=576/36 =16 hence x=4 hence the lengths of the sides are 68,40,36

OpenStudy (anonymous):

Thanks man. :d

OpenStudy (anonymous):

need not mention...